How to use ref and cite?

Marek Rychlik September 3, 2003

1 Equation

Let us try some equations:

A = πr^{2}  (1)

This is the second page. 2

Thist is the third page. Did you like equation (1)? Emacs commands generally involve the CONTROL key (sometimes labeled CTRL or CTL) or the META key (sometimes labeled EDIT or ALT). Rather than write that in full each time, we’ll use the following abbreviations:

C-¡chr¿ means hold the CONTROL key while typing the character ¡chr¿ Thus, C-f would be: hold the CONTROL key and type f. M-¡chr¿ means hold the META or EDIT or ALT key down while typing ¡chr¿. If there is no META, EDIT or ALT key, instead press and release the ESC key and then type ¡chr¿. We write ¡ESC¿ for the ESC key.

Important note: to end the Emacs session, type C-x C-c. (Two characters.) The characters ”¿¿” at the left margin indicate directions for you to try using a command. For instance:

The symbol˜ prevents a line break. The symbol p̃revents a line break.

See the difference between a¡b and a < b.

Unnumbered equation:

a < b

or

a < b

Numbered equation:

a < b  (2)